Please help ASAP with my algebra question. We are on quadratic equations and I am stuck on a word problem:
The area of a square can be doubled by increasing the length by 6 inches and the width by 4 inches. What is the length of the side of the square?

Answer by dfrazzetto(283)


square to L=W
original area = LxL = L^2
(L+6)(L+4) = 2l^2
L^2 + 10L + 24 = 2L^2
L^2 - 10L - 24 = 0
(L-12)(L+2) = 0
L = 12, -2
12 inches